# Interpretability-Guided Bi-objective Optimization: Aligning Accuracy and Explainability

### v1

@@ Abstract
This paper introduces Interpretability-Guided Bi-objective Optimization (IGBO), a framework that trains interpretable models by incorporating structured domain knowledge via a bi-objective formulation. IGBO encodes feature importance hierarchies as a Directed Acyclic Graph (DAG) and uses Temporal Integrated Gradients (TIG) to measure feature importance. To address the Out-of-Distribution (OOD) problem in TIG computation, we propose an Optimal Path Oracle that learns data-manifold-aware integration paths. Theoretical analysis proves convergence properties and robustness to mini-batch noise, while empirical results on time-series data demonstrate IGBO’s effectiveness in enforcing DAG constraints with minimal accuracy loss, outperforming standard regularization baselines.
Keywords: Interpretable Machine Learning, Bi-objective Optimization, Explainable AI (XAI), Temporal Integrated Gradients, Optimal Path Oracle, Structured Regularization
@@ I Introduction
To build more trustworthy systems, interpretability must be integrated directly into the training objective. A common approach adds a regularization term penalizing undesirable behavior. However, such methods often treat interpretability as a monolithic, soft constraint, offering practitioners limited control over the precise trade-off between accuracy and interpretability. More critically, they lack a formal mechanism to enforce structured, relational constraints that reflect complex domain expertise, such as hierarchies of feature importance. Additionally, gradient-based attribution methods like Integrated Gradients suffer from the Out-of-Distribution (OOD) problem when computing gradients along straight-line paths between inputs and baselines.
This work proposes a principled framework to address these gaps: Interpretability-Guided Bi-objective Optimization (IGBO). IGBO formalizes interpretable model training as two interconnected bi-objective optimization problems: (1) training the main model $F_{\theta}$ to simultaneously minimize a primary task loss $\mathcal{L}(\theta)$ and an interpretability loss $\mathcal{H}(\theta)$, and (2) training an Optimal Path Oracle $G$ to generate data-manifold-aware integration paths for robust gradient computation. The interpretability loss is derived from two key components: (1) a Temporal Integrated Gradients (TIG) metric that provides a differentiable measure of feature importance over time, and (2) a Directed Acyclic Graph (DAG) defined by a domain expert, which encodes permissible relative importance relationships between features (e.g., feature $A$ should be more important than feature $B$).
@@
Contributions. This paper makes the following contributions:
  • We introduce the IGBO framework, which integrates structured domain knowledge (via a DAG) into gradient-based learning through a differentiable TIG-based importance function $H_{k}(\mathbf{X},\theta)$, addressing both interpretability constraints and OOD issues in attribution.
@@
  • We establish a rigorous empirical validation protocol to assess IGBO’s capability to enforce complex DAG constraints on both synthetic and real-world time-series data, demonstrating its effectiveness with minimal compromise on predictive accuracy.
@@ II Related Work
Gradient-based Attribution and Robust Path Methods: Our Temporal Integrated Gradients (TIG) builds directly upon Integrated Gradients (IG) [sundararajan2017axiomatic], inheriting its axiomatic properties. Recent work has identified limitations in IG, particularly its sensitivity to the choice of baseline [sturmfels2020visualizing] and the fact that straight-line integration paths often traverse Out-of-Distribution (OOD) regions, leading to unreliable attributions [hess2020middle]. Several approaches attempt to address these issues: Expected Gradients [erion2019learning] averages over multiple baselines; BlurIG [xu2020explainable] applies smoothing; and adversarial training methods aim to produce more robust attributions [giove2023robust]. Our Optimal Path Oracle introduces a novel direction by learning a data-manifold-aware integration path, fundamentally changing how the attribution integral is computed. This approach is conceptually related to works exploring shortest paths on manifolds for explanations [garreau2020explaining] but differs in its fully differentiable, learned formulation and integration into a bi-objective training loop.
@@
Evaluation of Explanations and Benchmarking: Assessing explanation quality remains a core challenge in XAI due to the frequent lack of ground truth [mironicolau2024comprehensive]. Recent work proposes synthetic benchmarks [agarwal2023evaluating], axiomatic evaluation frameworks [hedstrom2022quantus], and methods to measure explanation robustness [yeh2019completeness]. Our use of an expert-defined DAG provides a concrete, testable prior for evaluating the alignment of learned feature importance with domain knowledge, contributing to research on measurable explanation fidelity. This approach aligns with recent calls for more rigorous, application-grounded evaluation of interpretability methods [doshi2017towards, lyon2022human].
@@ III Methodology
The Interpretability-Guided Bi-objective Optimization (IGBO) framework consists of two interconnected bi-objective optimization processes, both of which rely on gradient-based techniques for balancing competing objectives. These processes are:
@@
    Training an optimal path generator $G_{K}$ to produce data-manifold-aware integration paths for robust gradient computation in Temporal Integrated Gradients (TIG).
@@ III-A Interpretability Prior Definition
Sequential Model Definition.
@@
Temporal Integrated Gradients (TIG).
@@
This integral aggregates the sensitivity of the model’s output to feature $(t,k)$ along the entire path $\gamma$. The standard implementation uses the straight‑line path $\gamma_{\text{lin}}(\alpha)=\alpha\mathbf{X}+(1-\alpha)\mathbf{X}^{\prime}$. However, a major drawback of this linear interpolation is that intermediate points $\gamma_{\text{lin}}(\alpha)$ often lie outside the training data distribution (Out‑of‑Distribution, OOD). Since $F_{\theta}$ is not trained to behave consistently in OOD regions, the resulting gradients—and hence the attributions—can be unreliable and noisy.
III-A1 Routing Oracle: Learning Data‑Manifold‑Aware Integration Paths
Motivation and Design Philosophy.
To circumvent the OOD problem inherent in straight‑line integration paths, we propose a Routing Oracle (also referred to as a Pathfinder) that learns to generate piecewise‑linear paths adhering to the data manifold while minimizing total path length. By ensuring all intermediate integration points reside within regions where the model has been properly trained, the oracle yields stable and reliable gradient computations for TIG.
Flexible Implementation Options.
@@ III-A1 Routing Oracle: Learning Data‑Manifold‑Aware Integration Paths
  • Path complexity control: The number of anchor points $K$ can be tuned according to the available computational budget. Larger $K$ allows the path to better follow the data manifold but increases inference cost. For deterministic high‑throughput deployment, $K$ can also be fixed as a hyperparameter.
@@
Oracle Architecture and Recurrence Structure.
@@
$\mathbf{p}_{1},\dots,\mathbf{p}_{K}]=G(\mathbf{X}^{\prime},\mathbf{X},K;\theta_{G}).$
@@
where $r$ is a learned function. The explicit inclusion of the remaining‑step count (Ki+1)(K-i+1) allows the network to plan the path efficiently and adapt its behavior as the endpoint approaches.
Two‑Phase Training Procedure.
@@
    Discriminator pre‑training: We employ the discriminator $D(\cdot;\theta_{D})$ from a pre‑trained Generative Adversarial Network (GAN) as a frozen validity assessor. This discriminator, already trained to distinguish real data from synthetic samples, provides a reliable estimate of $\Pr(\mathbf{p}\in\mathcal{A})$ without requiring additional training.
@@
    $\mathcal{L}_{G}(\theta_{G})=\mathbb{E}_{(K,\mathbf{X}^{\prime},\mathbf{X})}\bigl[\mathcal{L}_{\text{path}}+\beta\cdot\mathcal{L}_{\text{valid}}\bigr],$ (4)
@@
    The hyper‑parameter $\beta>0$ controls the trade‑off between the two objectives.
Integration Strategy:
@@
  • Consistent evaluation: All models are assessed using identical integration paths, allowing fair comparison of their feature‑importance patterns.
@@
  • Streamlined deployment: New models can be interpreted immediately using a pre‑existing, validated oracle, which is especially valuable in regulated domains (e.g., healthcare or finance) where explanation consistency is critical.
III-B Feature Importance Function $H_{k}(\mathbf{X},\theta)$
To obtain a single, global importance score for feature $k$ across the entire sequence for a specific input, we aggregate the temporal attributions. We define the importance function as:
$H_{k}(\mathbf{X},\theta)=\frac{\sum_{i=1}^{T}\sum_{j=1}^{T}\left|\text{TIG}_{i,k}(\mathbf{X},\theta)[j]\right|}{T(T+1)/2}.$ (8)
The denominator normalizes the score by the total absolute deviation of feature $k$ from its baseline, rendering $H_{k}$ a measure of the model’s sensitivity per unit change in the input feature. This normalization ensures importance scores are comparable across features with different native scales or units. The function $H_{k}:\mathcal{X}\times\Theta\to\mathbb{R}_{\geq 0}$ is central to our framework, providing a differentiable mapping from the joint space of inputs and parameters to a non-negative importance score.
Interpretability Graph (DAG) as a Structured Prior: Domain expertise is encoded as a Directed Acyclic Graph $G=(V,E)$. Each node $v\in V$ corresponds to an input feature. A directed edge $u\rightarrow v$ encodes the semantic prior that, on average across the data distribution $\mathcal{D}$, feature $u$ should be more important than feature $v$. To allow for controlled flexibility and avoid overly rigid constraints, we formalize this as the requirement that the expected difference in importance scores lies within a specified interval:
@@ III-C Interpretability Loss
To integrate the DAG constraint from Eq. (9) into the training objective, we construct a differentiable interpretability loss function $\mathcal{H}:\Theta\to\mathbb{R}_{\geq 0}$. Since the expectation in Eq. (9) is over the data distribution, we work with its Monte Carlo estimate computed on a mini-batch.
@@
$\begin{split}\mathcal{H}(\theta;B)=\sum_{(u,v)\in E}\Big[&\max(\epsilon_{u,v}-d_{u,v}(\theta;B),\;0)\\ +&\max(d_{u,v}(\theta;B)-\delta_{u,v},\;0)\Big].\end{split}$ (11)
@@
III-D Gradient-Based Optimization via Projection Mapping

### v2

@@ Abstract
This paper introduces Interpretability-Guided Bi-objective Optimization (IGBO), a framework that trains interpretable models by incorporating structured domain knowledge via a bi-objective formulation. IGBO encodes feature importance hierarchies as a Directed Acyclic Graph (DAG) via Central Limit Theorem-based construction and uses Temporal Integrated Gradients (TIG) to measure feature importance. To address the Out-of-Distribution (OOD) problem in TIG computation, we propose an Optimal Path Oracle that learns data-manifold-aware integration paths. Theoretical analysis establishes convergence properties via a geometric projection mapping $\mathcal{P}$ and proves robustness to mini-batch noise. Central Limit Theorem-based construction of the interpretability DAG ensures statistical validity of edge orientation decisions. Empirical results on time-series data demonstrate IGBO’s effectiveness in enforcing DAG constraints with minimal accuracy loss, outperforming standard regularization baselines.
Keywords: Interpretable Machine Learning, Bi-objective Optimization, Explainable AI (XAI), Temporal Integrated Gradients, Optimal Path Oracle, Structured Regularization, Central Limit Theorem
@@ I Introduction
To build more trustworthy systems, interpretability must be integrated directly into the training objective. A common approach adds a regularization term penalizing undesirable behavior. However, such methods often treat interpretability as a monolithic, soft constraint, offering practitioners limited control over the precise trade-off between accuracy and interpretability. More critically, they lack a formal mechanism to enforce structured, relational constraints that reflect complex domain expertise, such as hierarchies of feature importance. Additionally, gradient-based attribution methods like Integrated Gradients suffer from the OOD problem when computing gradients along straight-line paths between inputs and baselines.
This work proposes a principled framework to address these gaps: Interpretability-Guided Bi-objective Optimization (IGBO). IGBO formalizes interpretable model training as two interconnected bi-objective optimization problems: (1) training the main model $F_{\theta}$ to simultaneously minimize a primary task loss $\mathcal{L}(\theta)$ and an interpretability loss $\mathcal{H}(\theta)$ via a geometric projection mapping $\mathcal{P}$ (Algorithm 1), and (2) training an Optimal Path Oracle $G$ to generate data-manifold-aware integration paths for robust gradient computation. The interpretability loss is derived from two key components: (1) a Temporal Integrated Gradients (TIG) metric that provides a differentiable measure of feature importance over time, and (2) a Directed Acyclic Graph (DAG) defined via Central Limit Theorem-based construction (Section III-D) that encodes permissible relative importance relationships between features (e.g., feature $A$ should be more important than feature $B$).
@@
Application: The IGBO framework not only provides immediate interpretability guarantees but also opens avenues for human-in-the-loop model alignment and task-specific model specialization, which we explore in later sections (Section III, Discussion).
Contributions: This paper makes the following contributions:
  • We introduce the IGBO framework, which integrates structured domain knowledge (via a DAG) into gradient-based learning through a differentiable TIG-based Argument Satisfaction function $H_{k}(\mathbf{X},\theta)$, addressing both interpretability constraints and OOD issues in attribution.
@@
  • We establish theoretical guarantees including convergence to Pareto-stationary points via projection mapping $\mathcal{P}$ (Theorems V-B3, V-B4), robustness to gradient noise (Theorem V-C1), and statistical validity of DAG construction via Central Limit Theorem approximation (Section III-D).
  • We provide empirical validation demonstrating IGBO’s capability to enforce complex DAG constraints on both synthetic and real-world time-series data with minimal compromise on predictive accuracy.
@@ II Related Work
Gradient-based Attribution and Robust Path Methods: Our TIG builds directly upon Integrated Gradients (IG) [sundararajan2017axiomatic], inheriting its axiomatic properties. Recent work has identified limitations in IG, particularly its sensitivity to the choice of baseline [sturmfels2020visualizing] and the fact that straight-line integration paths often traverse OOD regions, leading to unreliable attributions [hess2020middle]. Several approaches attempt to address these issues: Expected Gradients [erion2019learning] averages over multiple baselines; Blurring Integrated Gradients [xu2020explainable] applies smoothing; and adversarial training methods aim to produce more robust attributions [giove2023robust]. Our Optimal Path Oracle introduces a novel direction by learning a data-manifold-aware integration path, fundamentally changing how the attribution integral is computed. This approach is conceptually related to works exploring shortest paths on manifolds for explanations [garreau2020explaining] but differs in its fully differentiable, learned formulation and integration into a bi-objective training loop.
@@
Probabilistic Graph Construction: Our method for constructing the interpretability DAG via Central Limit Theorem approximation (Section III-D) relates to statistical methods for causal discovery [spirtes2000causation] and confidence-based edge orientation [kalisch2007estimating]. However, unlike causal discovery methods that infer structure from data, our approach uses statistical confidence to operationalize expert knowledge about feature importance hierarchies, with theoretical guarantees on acyclicity (Corollary V-A6) and transitivity (Theorem V-A5).
@@ III Methodology
The IGBO framework consists of two interconnected bi-objective optimization processes, both of which rely on gradient-based techniques for balancing competing objectives. These processes are:
@@
    Training an optimal path generator $G$ to produce data-manifold-aware integration paths for robust gradient computation in TIG.
@@ III-A Interpretability Prior Definition
Sequential Model Definition
@@
Temporal Integrated Gradients.
@@
This integral aggregates the sensitivity of the model’s output to feature $(t,k)$ along the entire path $\gamma$. The standard implementation uses the straight-line path $\gamma_{\text{lin}}(\alpha)=\alpha\mathbf{X}+(1-\alpha)\mathbf{X}^{\prime}$. However, a major drawback of this linear interpolation is that intermediate points $\gamma_{\text{lin}}(\alpha)$ often lie outside the training data distribution (OOD). Since $F_{\theta}$ is not trained to behave consistently in OOD regions, the resulting gradients—and hence the attributions—can be unreliable and noisy.
III-A1 Routing Oracle: Learning Data-Manifold-Aware Integration Paths
Motivation and Design Philosophy
To circumvent the OOD problem inherent in straight-line integration paths, we propose a Routing Oracle (also referred to as a Pathfinder) that learns to generate piecewise-linear paths adhering to the data manifold while minimizing total path length. By ensuring all intermediate integration points reside within regions where the model has been properly trained, the oracle yields stable and reliable gradient computations for TIG.
Flexible Implementation Options
@@ III-A1 Routing Oracle: Learning Data-Manifold-Aware Integration Paths
  • Path complexity control: The number of anchor points $K$ can be tuned according to the available computational budget. Larger $K$ allows the path to better follow the data manifold but increases inference cost. For deterministic high-throughput deployment, $K$ can also be fixed as a hyperparameter.
@@
Oracle Architecture and Recurrence Structure
@@
$[\mathbf{p}_{1},\dots,\mathbf{p}_{K}]=G(\mathbf{X}^{\prime},\mathbf{X},K;\theta_{G}).$
@@
where $r$ is a learned function. The explicit inclusion of the remaining-step count (Ki+1)(K-i+1) allows the network to plan the path efficiently and adapt its behavior as the endpoint approaches.
Two-Phase Training Procedure
@@
    Discriminator pre-training: We employ the discriminator $D(\cdot;\theta_{D})$ from a pre-trained Generative Adversarial Network (GAN) as a frozen validity assessor. This discriminator, already trained to distinguish real data from synthetic samples, provides a reliable estimate of $\Pr(\mathbf{p}\in\mathcal{A})$ without requiring additional training.
@@
    $\mathcal{L}_{G}(\theta_{G})=\mathbb{E}_{(K,\mathbf{X}^{\prime},\mathbf{X})}\bigl[\mathcal{L}_{\text{path}},\,\mathcal{L}_{\text{valid}}\bigr],$ (4)
@@
    Training the oracle is a bi-objective optimization problem.
Integration Strategy
@@
  • Consistent evaluation: All models are assessed using identical integration paths, allowing fair comparison of their feature-importance patterns.
@@
  • Streamlined deployment: New models can be interpreted immediately using a pre-existing, validated oracle, which is especially valuable in regulated domains (e.g., healthcare or finance) where explanation consistency is critical.
III-B Argument Satisfaction Function $H_{k}(\mathbf{X},\theta)$
We aggregate temporal attributions to obtain a single scalar score for each feature $k$ over the entire input sequence. Rather than representing raw feature importance, $H_{k}(\mathbf{X},\theta)$ is interpreted as an Argument Satisfaction score: it quantifies how well the cumulative Temporal Integrated Gradients of feature $k$ explain the model’s deviation from a chosen baseline across time.
At each time step, we measure the residual between the observed output change and the portion attributed to feature $k$, and convert this residual into a bounded agreement score via a sigmoid mapping:
$\begin{split}h_{t,k}(\mathbf{X},\theta)&=\left|F_{\theta}(\mathbf{X})[t]-F_{\theta}(\mathbf{X}^{\prime})[t]-\sum_{i=1}^{t}\text{TIG}_{i,k}[t]\right|,\\ H_{k}(\mathbf{X},\theta)&=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{1+\beta\cdot h_{t,k}}\in[0,1],\;\beta>0.\end{split}$ (8)
Higher values of $H_{k}$ indicate stronger satisfaction of the explanatory argument that feature $k$ contributes meaningfully to the model’s prediction. The normalization ensures comparability across features and yields a smooth, differentiable quantity suitable for optimization.
Interpretability Graph DAG as a Structured Prior: Domain expertise is encoded as a Directed Acyclic Graph $G=(V,E)$. Each node $v\in V$ corresponds to an input feature. A directed edge $v\rightarrow u$ implies the semantic prior that, on average over the data distribution $\mathcal{D}$, feature $u$ should exhibit a higher Argument Satisfaction score than feature $v$. To allow for controlled flexibility and avoid overly rigid constraints, we formalize this as the requirement that the expected difference in importance scores lies within a specified interval:
@@ III-C Interpretability Loss
To integrate the DAG constraint from Eq. (9) into the training objective, we construct a differentiable interpretability loss function $\mathcal{H}:\Theta\to[0,1]$. Since the expectation in Eq. (9) is over the data distribution, we work with its Monte Carlo estimate computed on a mini-batch.
@@
$\begin{split}\mathcal{H}(\theta;B)=\frac{1}{|E|}\sum_{(u,v)\in E}\Big[&\max(\epsilon_{u,v}-d_{u,v}(\theta;B),\;0)\\ +&\max(d_{u,v}(\theta;B)-\delta_{u,v},\;0)\Big].\end{split}$ (11)
@@
III-D Graph Construction via Central Limit Approximation
This subsection gives a concise, implementation-oriented description of the DAG construction procedure and points to the formal probabilistic statements and proofs collected in Section V-A.
Overview
We build a directed edge $u\to v$ when the batch-mean importance score of feature $u$ is sufficiently larger than that of $v$ with high confidence. All probabilistic claims below are justified by the CLT-based lemmas and theorems in Section V-A.
Per-batch statistics
For a batch $B$ of size $|B|$ define the batch mean
$s_{k}(B)=\frac{1}{|B|}\sum_{\mathbf{X}\in B}H_{k}(\mathbf{X},\theta).$
Under the assumptions of Lemma V-A2 (i.i.d. sampling and finite second moments) each $s_{k}(B)$ is approximately normal with mean $\mu_{k}$ and variance $\mathrm{Var}(H_{k})/|B|$. When multiple independent batches $B^{(1)},\dots,B^{(N)}$ are available, the per-batch means $s_{k}^{(i)}$ are i.i.d.; see Lemma V-A2 and the discussion that follows it.
Edge orientation rule (practical form)
For a candidate ordered pair $(u,v)$ let $d_{u,v}^{(i)}=s_{u}^{(i)}-s_{v}^{(i)}$. Using the CLT and the variance expression of Lemma V-A3 we obtain the normal approximation
$d_{u,v}^{(i)}\approx\mathcal{N}\!\Big(\mu_{u}-\mu_{v},\;\sigma_{uv}^{2}\Big),$
$\sigma_{uv}^{2}=\frac{\mathrm{Var}(H_{u})+\mathrm{Var}(H_{v})-2\mathrm{Cov}(H_{u},H_{v})}{|B|}.$
Hence the plug-in estimator for the edge probability is
$\widehat{P}(s_{u}>s_{v})=\Phi\!\Big(\frac{\bar{s}_{u}-\bar{s}_{v}}{\sqrt{\widehat{\mathrm{Var}}(s_{u}-s_{v})}}\Big),$
where $\widehat{\mathrm{Var}}(s_{u}-s_{v})$ is obtained either by within-batch variance divided by $|B|$ (Lemma V-A3) or by the empirical between-batch variance of $\{s_{u}^{(i)}-s_{v}^{(i)}\}_{i=1}^{N}$. The numerical decision is
$u\to v\quad\Longleftrightarrow\quad\widehat{P}(s_{u}>s_{v})>\alpha,$
for a chosen $\alpha>0.5$. The equivalence between the probabilistic rule and the mean-ordering (for $\alpha=0.5$) is formalized in Theorem V-A4.
Acyclicity and transitivity
- If $\alpha=0.5$, Theorem V-A4 shows the rule reduces to ordering by population means $\mu_{k}$; orienting edges by this total order yields a DAG (Corollary V-A6). - For $\alpha>0.5$, transitivity (and hence acyclicity) follows under the margin/variance conditions made explicit in Theorem V-A5; the required triangle‑type bound on standard deviations is a direct consequence of the $L^{2}$ triangle inequality (see the proof of Theorem V-A5).
Practical recommendations (summary)
Refer to Section V-A for formal statements and proofs. In practice:
  • Prefer the between-batch empirical variance of $s_{k}$ when $N$ is moderate to large (robust uncertainty quantification).
  • If only a single batch is available, estimate $\mathrm{Var}(H_{k})$ with the unbiased sample variance (divide by |B|1|B|-1) and obtain $\widehat{\mathrm{Var}}(s_{k})=\widehat{\mathrm{Var}}(H_{k})/|B|$.
  • Use bootstrap or permutation tests and/or truncated‑normal corrections when $|B|$ is small or $\mu_{k}$ is near the boundaries $[0,1]$ (see Remark V-A7).
  • To ensure transitivity in practice, check the margin/variance conditions of Theorem V-A5 or increase $|B|$ / $\alpha$ as needed.
III-E Gradient-Based Optimization via Projection Mapping